OpenStudy (kitkat1):

Given that f(x) = 2x −5, find the value of x that makes f(x) = 15. 5 10 20 25

OpenStudy (idealist10):

2x-5=15 2x=15+5 2x=20 x=20/2 x=10 Answer: 10

OpenStudy (texaschic101):

f(x) = 2(10) - 5 f(5) = 20 - 5 f(5) = 15 yep ^^ answer is 10

OpenStudy (kitkat1):

could u help me with another one?